# Twist, splay, and uniform domains in ferroelectric nematic liquid crystals

### v1

@@ II.3 Radial patterns
Figure 3: Structure of a +1 radial prepatterned splay defect. (a) A polarizing optical microscopy texture (recorded in a monochromatic light using a green filter with $532~{}\mathrm{nm}$ wavelength and $1~{}\mathrm{nm}$ bandwidth) of an $\mathrm{N}_{\mathrm{F}}$ in a cell of the thickness $h=5.5~{}\mu\mathrm{m}$ (DIO, temperature $T=60^{\circ}\mathrm{C}$). (b) Fitting the dependence of transmitted light intensity on the angle $\gamma$ between the polarizers yields the twist angle $\tau\approx 180^{\circ}$ for samples of different thickness in the range $3~{}\mu\mathrm{m}<h<16~{}\mu\mathrm{m}$.
Figure 4: A polarizing optical microscopy texture (recorded in a monochromatic light using a green filter with $532~{}\mathrm{nm}$ wavelength and $1~{}\mathrm{nm}$ bandwidth) of the $\mathrm{N}_{\mathrm{F}}$ phase of (a) pure DIO, $T=50^{\circ}\mathrm{C}$ and (b) DIO doped with BMIM-$\mathrm{PF}_{\mathrm{6}}$, $T=43^{\circ}\mathrm{C}$ in cells with radial patterns of the thickness $h=1.9~{}\mu\mathrm{m}$. (c) Number of domains versus distance to the +1 defect center for pure DIO and DIO doped with BMIM-$\mathrm{PF}_{\mathrm{6}}$.
@@
The radial pattern of the +1 defect in the N and $\mathrm{SmZ}_{\mathrm{A}}$ phases demonstrate smooth splay deformation of the director. In the $\mathrm{N}_{\mathrm{F}}$ phase, the textures show domains of two types. The first are “pie-slices”, or splay domains (SDs), with $\mathbf{P}$ parallel or anti-parallel to the radial direction $\hat{\mathbf{r}}$, pointing either away from the core of the +1 defect at $\hat{\mathbf{r}}=0$ or towards it, as shown in Fig. 2(a,b). As established by polarizing microscopy, within each SD, $\mathbf{P}$ does not twist along the $z$ axis, except perhaps within the domain walls that separate splay sectors of antiparallel $\mathbf{P}$. These domains are similar to the UDs in the planar cells in the sense that the polarization does not twist. The second type are $\pi$-TDs, similar to the $\pi$-TDs in the planar cells: $\mathbf{P}$ twists around the $z$ axis by $\tau\approx 180^{\circ}$, as shown in Fig. 3(a,b).
@@
The frequency of domain appearance depends on the cell thickness and temperature. Thick cells, $h>6~{}\mathrm{\mu m}$, upon cooling below $61^{\circ}\mathrm{C}$, show exclusively TDs, Fig. 2(e,g). Cells of an intermediate thickness, $1~{}\mathrm{\mu m}<h<6~{}\mathrm{\mu m}$, show both $\pi$-TDs and SDs, Fig. 2(e,g). Structures in cells with $h\leq 1~{}\mathrm{\mu m}$ are either uniformly radial, Fig. 2(h), or show a few SDs. In thin cells, $1~{}\mathrm{\mu m}<h<6~{}\mathrm{\mu m}$, the domains form at about $(5-10)^{\circ}\mathrm{C}$ below the $\mathrm{SmZ}_{\mathrm{A}}$-$\mathrm{N}_{\mathrm{F}}$ transition point. The thickness dependency of the temperature at which the domains appear can be qualitatively explained by the temperature dependence of the polarization magnitude $P_{0}$ which increases from about $P_{0}=3.4\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$ to $P_{0}=4.6\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$ Nishikawa et al. (2017a) as the temperature decreases following the $\mathrm{SmZ}_{\mathrm{A}}$-$\mathrm{N}_{\mathrm{F}}$ transition; the bound charge effects leading to the domains might be weaker than the stabilizing surface anchoring and ion screening.
@@
Within the range of coexistence, the fraction of the SD domains increases as the cell thickness $h$ decreases, Fig. 2(e,g). The number of domains increases with the cooling rate, Fig. 2(c,d), which is natural as the higher cooling rate does not leave time for the domains to coalesce and bring the system closer to the equilibrium. Another potential reason is that for a longer cooling time, the highly polar material absorbs ions from the surroundings, such as glue, BY layer, etc., to screen the bound charge $\rho=-\nabla\cdot\mathbf{P}$. The ion concentration of free ions in the $\mathrm{N}_{\mathrm{F}}$ cannot be measured directly by conventional techniques since the polarization reorientation reduces the electric field in the $\mathrm{N}_{\mathrm{F}}$ bulk to zero Clark et al. (2024). Some support of the idea that the concentration of ions can increase with time is provided by measurement in the N phase, by briefly heating the sample to $120^{\circ}\mathrm{C}$; the concentration is found to increase from $c(0)=5.0\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$ at the start of experiment to $c(18~{}\mathrm{hours})=6.3\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$ after 18 hours of keeping the sample in the $\mathrm{N}_{\mathrm{F}}$ phase at $65^{\circ}\mathrm{C}$.
The number of domains decreases significantly when DIO is doped with the ionic fluid BMIM-$\mathrm{PF}_{\mathrm{6}}$, Fig. 4. At a weight concentration 0.5 wt % of BMIM-$\mathrm{PF}_{\mathrm{6}}$, which corresponds to the concentration of ions $1.5\times 10^{25}~{}\mathrm{ions}/\mathrm{m}^{3}$ when added molecules are fully ionized, the $\mathrm{N}_{\mathrm{F}}$ phase preserves its ferroelectric ordering, as reported by Zhong et al. Zhong et al. (2025). The free ions screen the space charges created by the splay of polarization and help to minimize the electrostatic energy.
III Spontaneous $\mathbf{P}$ twist
To begin the theoretical analysis, it is worth to first review Khachaturyan’s theoretical prediction from 1975 Khachaturyan (1975) that $\mathrm{N}_{\mathrm{F}}$s in an infinite sample (thick film) may spontaneously develop twisted polarization $\mathbf{P}$ domains of linear size $R$ with a characteristic wavelength $\lambda_{z}$. Experiments on thin films Kumari et al. (2024a) provide qualitative support of this prediction.
III.1 Twist in a cylinder
@@
We will consider a cylindrical domain with radius $R$ and length $h$, as shown in Fig. 5. The intuitive explanation is that, in the absence of twist, a uniform polarization ($\mathbf{P}=P_{0}\hat{\mathbf{x}}$, say) will generate a strong depolarization field due to the accumulation of uncompensated charge on the domain boundary, as shown on the left panel of Fig. 5. The charges can be partially compensated by twisting the direction of $\mathbf{P}$ along the cylinder, as demonstrated on the right panel of Fig. 5. We can calculate the optimal period $\lambda_{z}$ of the twist by balancing this electrostatic energy gain with the elastic cost of the deformation.
@@
Figure 5: A cylindrical domain with uniform polarization $\mathbf{P}=P_{0}\hat{\mathbf{x}}$ (left panel) incurs an energetic cost due to the uncompensated charges and consequent depolarization field $\mathbf{E}_{\mathrm{dep}}$. Instead, the polarization $\mathbf{P}$ can twist (right panel) with a certain period $\lambda_{z}$ and create regions of alternating charge, thereby partially mitigating the energetic cost of the depolarization field. This twist is balanced by the elastic cost of the twist.
@@
To be more specific, the $\mathrm{N}_{\mathrm{F}}$ has a continuously varying polarization vector $\mathbf{P}\equiv\mathbf{P}(\mathbf{r})=P_{0}\mathbf{n}(\mathbf{r})$ within the material, with some fixed magnitude $|\mathbf{P}|=P_{0}$ and a varying orientation $\mathbf{n}(\mathbf{r})$. Spatial variations in $\mathbf{P}$ or domain boundaries will generate a bound charge distribution given by $\rho=-\nabla\cdot\mathbf{P}$. This distribution of charge is energetically costly and we can calculate the corresponding (screened) electrostatic energy as
@@
where $\epsilon$ is the relative dielectric constant of the material and $\kappa\equiv 1/\lambda_{D}$ is the (inverse) Debye screening length. The screening comes from the free ions present in the material, with $\kappa\approx\sqrt{n}\,e/\sqrt{\epsilon\epsilon_{0}k_{B}T}$ for monovalent ions with concentration $n$, with $e$ the fundamental charge, $\epsilon\epsilon_{0}$ the material permittivity, and $k_{B}T$ the thermal energy. This free energy $F_{\rho}$ can be expressed in Fourier space as
@@
where $\tilde{\mathbf{P}}_{\mathbf{k}}\equiv\int\mathrm{d}\mathbf{r}\,e^{-i\mathbf{k}\cdot\mathbf{r}}\mathbf{P}(\mathbf{r})$. Note that even in cases where $\rho(\mathbf{r})$ vanishes in the bulk of a sample or domain, we may still get contributions to $F_{\rho}$ due to uncompensated charges at the boundaries (the depolarization field).
@@
$F$ =${}_{n}=\int\mathrm{d}\mathbf{r}\left[\frac{K_{1}}{2}(\nabla\cdot\mathbf{n})^{2}+\frac{K_{2}}{2}(\mathbf{n}\cdot(\nabla\times\mathbf{n}))^{2}\right.$
$\qquad\qquad\qquad\left.{}+\frac{K_{3}}{2}(\mathbf{n}\times(\nabla\times\mathbf{n}))^{2}\right],$ (3)
with $K_{1}$, $K_{2}$, and $K_{3}$ the splay, twist, and bend elastic constants, respectively de Gennes and Prost (1995). To see how $F_{n}$ and $F_{\rho}$ compete to create variations in $\mathbf{P}$, we look for free energy minimizers which only vary along the vertical direction $z$:
@@
$\sum_{m=-\infty}^{\infty}A_{m}A^{*}_{m-n}=\begin{cases}1&n=0\\ 0&n\neq 0\end{cases}.$ (7)
Provided we have a thick sample with $h\gg\lambda_{D}$, we substitute Eqs. (4,6) into Eq. (2) and find that
@@
where $\alpha_{n}\equiv\,R\sqrt{[k_{z}(n+1)]^{2}+\kappa^{2}}$ and $I_{1}(\alpha)$, $K_{1}(\alpha)$ are the modified Bessel functions of the first and second kind, respectively. It is worth noting that $I_{1}(\alpha)K_{1}(\alpha)$ is a monotonically decreasing function of $\alpha$, meaning that the electrostatic energy favors large values of $\alpha_{n}\sim k_{z}$. As we shall see, the elastic contribution will favor small values of $k_{z}$, instead.
@@
which clearly is minimized for $k_{z}\rightarrow 0$. We can also see that higher order modes $A_{n}$ with $|n|>0$ always cost more elastic energy. This is less obvious for the electrostatic interaction in Eq. (8), but it is possible to argue on more general grounds Khachaturyan (1975) that there is a stable free energy minimum with $A_{n}=0$ for all $|n|>0$.
@@
where $\alpha_{0}=R\sqrt{k_{z}^{2}+\kappa^{2}}$ and we recognize that with just the single mode we must have $\sum_{n}|A_{n}|^{2}=A_{0}^{2}=(A_{0}^{*})^{2}=1$ from Eq. (7). The total free energy in Eq. (11) now can be minimized with respect to $k_{z}$. We look for solutions for large domain size $R$ such that $R/\lambda_{D}\gg 1$. In this case, $\alpha_{0}\gg 1$ and we can make use of the asymptotic expansion $I_{1}(\alpha_{0})K_{1}(\alpha_{0})\approx(2\alpha_{0})^{-1}$. We find a minimum free energy at $k_{z}=k_{z}^{*}$, which corresponds to a preferred pitch $\lambda_{z}=2\pi/k_{z}^{*}$ of
@@
Substituting in reasonable values $\epsilon=100$, $P_{0}\approx 4.4\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$, $K_{2}=5~{}\mathrm{pN}$, $R=50~{}\mu\mathrm{m}$, and $T=350~{}\mathrm{K}$, and assuming no screening ($\kappa=0$) we find a pitch of $\lambda_{z}^{*}\approx 0.5~{}\mu\mathrm{m}$. This result is consistent with the previously reported data Kumari et al. (2024a) and with Fig. 1.
The pitch in Eq. (12) is analogous to the result derived by Khachaturyan Khachaturyan (1975), with some key corrections involving the contributions of screening. This result, along with the correction, was also recently found by Paik and Selinger using a different analysis Paik and Selinger (2024). We see that we get the twisted structure only when the screening is sufficiently weak: $\lambda_{D}^{2}>(4K_{2}R\epsilon\epsilon_{0})^{2/3}P_{0}^{-4/3}$. Assuming that the screening comes from some concentration $c$ of monovalent ions, then we have the squared Debye screening length $\lambda_{D}^{2}=\kappa^{-2}=\epsilon\epsilon_{0}k_{B}T/(ce^{2})$. The twisted state occurs only when the concentration of ions is below some critical value:

### v2

@@ Results
Figure 3: Twist in a +1 radial defect pre-patterned cell. a A polarizing optical microscopy texture (recorded in a monochromatic light using a green filter with $532~{}\mathrm{nm}$ wavelength and $1~{}\mathrm{nm}$ bandwidth) of an $\mathrm{N}_{\mathrm{F}}$ in a cell of the thickness $h=5.5~{}\mu\mathrm{m}$ (DIO, temperature $T=60^{\circ}\mathrm{C}$). The curved white arrow shows the direction of polarization twist and the anti-parallel arrows indicate the $\pi$-twist of $\mathbf{P}$ between the top and bottom plates. b Schematic of the polarization $\mathbf{P}$ (green arrow) performing a $\pi$-twist between the bottom and top cell surfaces (orange). c Fitting the dependence of transmitted light intensity [through a small region in a domain like the red rectangle in a] on the angle $\gamma$ between the polarizers yields the twist angle $\tau\approx 180^{\circ}$ for samples of different thickness in the range $3~{}\mu\mathrm{m}<h<8~{}\mu\mathrm{m}$, demonstrating the $\pi$-twist (see Methods). The error bars are the standard deviations of light intensity fluctuations in the measured small regions.
@@
The radial patterns of the +1 defect in the N and $\mathrm{SmZ}_{\mathrm{A}}$ phases demonstrate smooth splay deformation of the director. In the $\mathrm{N}_{\mathrm{F}}$ phase, the textures show domains of two types: The first are pie slices, or splay domains (SDs), with $\mathbf{P}$ parallel or anti-parallel to the radial direction $\hat{\mathbf{r}}$, pointing either away from the core of the +1 defect at $\hat{\mathbf{r}}=0$ or towards it, as shown in Fig. 2a,b. As established by polarizing microscopy, within each SD, $\mathbf{P}$ does not twist along the $z$ axis, except perhaps within the domain walls that separate sectors of antiparallel $\mathbf{P}$, similar to the UDs in the planar cells. The second type are $\pi$-TDs, similar to the $\pi$-TDs in the planar cells: $\mathbf{P}$ twists around the $z$ axis by $\tau\approx 180^{\circ}$, as shown in Fig. 3.
Thick cells, $h>6~{}\mathrm{\mu m}$, upon cooling below $61^{\circ}\mathrm{C}$, show exclusively $\pi$-TDs, Fig. 2e,g. Cells of an intermediate thickness, $1~{}\mathrm{\mu m}<h<6~{}\mathrm{\mu m}$, show both $\pi$-TDs and SDs, Fig. 2e,g. Structures in cells with $h\leq 1~{}\mathrm{\mu m}$ are either uniformly radial, Fig. 2h, or show a few SDs. In all of these cases, there is splay deformation originating from the $+1$ central vortex and the SDs and $\pi$-TDs terminate in a complex manner near this central defect region (i.e., at the tips of the pie slices), as seen in Fig. 2b. Thus, when comparing these samples to our theoretical results below, we will focus on the region away from this complex region, out at distances $R\approx 100~{}\mu\mathrm{m}$ from the defect center where the sectors do not terminate but where we still find substantial splay deformation and, consequently, the presence of bound charge.
Figure 4: Effects of ion addition. A polarizing optical microscopy texture (recorded in a monochromatic light using a green filter with $532~{}\mathrm{nm}$ wavelength and $1~{}\mathrm{nm}$ bandwidth) of the $\mathrm{N}_{\mathrm{F}}$ phase of, a, pure DIO, $T=50^{\circ}\mathrm{C}$ and b, DIO doped with 0.5% BMIM-$\mathrm{PF}_{\mathrm{6}}$, $T=43^{\circ}\mathrm{C}$ in cells with radial patterns of the thickness $h=1.9~{}\mu\mathrm{m}$. c Number of domains versus distance to the +1 defect center for pure DIO and DIO doped with BMIM-$\mathrm{PF}_{\mathrm{6}}$. Error bars indicate that it was difficult to distinguish about two of the sectors in the two samples.
@@
In thin cells, $1~{}\mathrm{\mu m}<h<6~{}\mathrm{\mu m}$, the domains form at about $(5-10)^{\circ}\mathrm{C}$ below the $\mathrm{SmZ}_{\mathrm{A}}$-$\mathrm{N}_{\mathrm{F}}$ transition point. The thickness dependency of the temperature at which the domains appear can be qualitatively explained by the temperature dependence of the polarization magnitude $P_{0}$ which increases from about $P_{0}=3.4\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$ to $P_{0}=4.6\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$ [2] as the temperature decreases following the $\mathrm{SmZ}_{\mathrm{A}}$-$\mathrm{N}_{\mathrm{F}}$ transition; the bound charge effects leading to the domains might be weaker than the stabilizing surface anchoring and ion screening.
The addition of either a $\pi$-TD or SD domain increases the effective charge of the central defect as the polarization orientation will twist by an additional factor of $2\pi$ for each pie slice in the domain configuration. The central $+1$ defect in the pattern, then, breaks up into a complex arrangement where all of the pie slice corners merge together. When there are no sectors, such as in thin cells with $h\sim 1~{}\mu\mathrm{m}$, the $+1$ defect splits into two $+1/2$ defect cores separated by a domain wall, as illustrated in Fig. 2h with a thin blue line of a somewhat smaller retardance ($\sim 160~{}\mathrm{nm}$ versus $\sim 210~{}\mathrm{nm}$ in the far-field). This result indicates that there is no melting of the ferroelectric order (in which case the domain wall will be absent) nor significant escape into the third dimension (in which case the central part would have significantly lower retardance). In a regular nematic $\mathrm{N}$, the $+1$ defect in the center would break up into disclination lines or contain an escaped structure [16]. The absence of the escape into the third dimension in the $\mathrm{N}_{\mathrm{F}}$ is also evident in other textures in Figs. 2, 3 and 4. This is not surprising since, in an $\mathrm{N}_{\mathrm{F}}$, such an escape of $\mathbf{P}$ along the normal to the cell would either deposit surface charges at the bounding plates or create additional splay with an accompanying bulk space charge. Note that even in a paraelectric N, a $+1$ defect prefers to split into a pair of $+1/2$ defects when the cell is sufficiently thin (see [16, 17, 18]).
@@
Within the range of coexistence, the fraction of the SD domains increases as the cell thickness $h$ decreases, Fig. 2e,g. The number of domains increases with the cooling rate, Fig. 2c,d. One potential reason is that for a longer cooling time, the highly polar material absorbs ions from the surroundings, such as glue, BY layer, etc., to screen the bound charge. The ion concentration of free ions in the $\mathrm{N}_{\mathrm{F}}$ cannot be measured directly by conventional techniques since polarization reorientation reduces the electric field in the $\mathrm{N}_{\mathrm{F}}$ bulk [19]. Briefly heating the sample to $120^{\circ}\mathrm{C}$ into the N phase and measuring the concentration supports the idea that ion concentration can increase in time: We find $c(0)=5.0\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$ at the start of experiment to $c(18~{}\mathrm{hours})=6.3\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$ after 18 hours of keeping the sample in the $\mathrm{N}_{\mathrm{F}}$ phase at $65^{\circ}\mathrm{C}$. The number of domains decreases significantly when DIO is doped with the ionic fluid BMIM-$\mathrm{PF}_{\mathrm{6}}$, Fig. 4. At a weight concentration 0.5 wt % of BMIM-$\mathrm{PF}_{\mathrm{6}}$, (an ion concentration of $1.5\times 10^{25}~{}\mathrm{ions}/\mathrm{m}^{3}$ for fully ionized added molecules), the $\mathrm{N}_{\mathrm{F}}$ phase preserves ferroelectric ordering, as reported by Zhong et al. [13].
Theory of twisted cylindrical domains
@@
To begin the theoretical analysis, we first review Khachaturyan’s theoretical prediction from 1975 [20] that $\mathrm{N}_{\mathrm{F}}$s may spontaneously develop twisted polarization $\mathbf{P}$ domains with a certain period $\lambda_{z}$. We consider a cylindrical domain with radius $R$ and length $h$, as shown in Fig. 5. In the absence of twist, a uniform polarization ($\mathbf{P}=P_{0}\hat{\mathbf{x}}$, say) will generate a strong depolarization field due to the accumulation of uncompensated charge on the domain boundary, as shown on the left panel of Fig. 5. The charges can be partially compensated by twisting the direction of $\mathbf{P}$ along the cylinder, as demonstrated on the right panel of Fig. 5. Note that this model is an idealization, as the polarization $\mathbf{P}$ in a real sample will typically not terminate with any component normal to the boundary, precisely to avoid the accumulation of the bound charge. There will thus likely be a complex deformation of the polarization near such a boundary. Although such configurations are beyond the scope of our analysis, we can use this idealized model to demonstrate the competition between electrostatics and elasticity in a simple way.
@@
Figure 5: Cylindrical domain schematic. A uniform polarization $\mathbf{P}=P_{0}\hat{\mathbf{x}}$ (left panel) incurs an energetic cost due to the uncompensated charges and consequent depolarization field $\mathbf{E}_{\mathrm{dep}}$. Instead, the polarization $\mathbf{P}$ can twist (right panel) with a certain period $\lambda_{z}$ and create regions of alternating charge, thereby partially mitigating the energetic cost of the depolarization field. This twist is balanced by the elastic cost of the twist.
@@
The $\mathrm{N}_{\mathrm{F}}$ has a continuously varying polarization vector $\mathbf{P}\equiv\mathbf{P}(\mathbf{r})=P_{0}\mathbf{n}(\mathbf{r})$ within the material, with some fixed magnitude $|\mathbf{P}|=P_{0}$ and a varying orientation $\mathbf{n}(\mathbf{r})$. The bound charge distribution is given by $\rho=-\nabla\cdot\mathbf{P}$. We can calculate the corresponding (screened) electrostatic energy as
@@
where $\epsilon$ is the relative dielectric constant of the material and $\kappa\equiv 1/\lambda_{D}$ is the (inverse) Debye screening length as derived in the Debye-Hückel approximation. We have $\kappa\approx\sqrt{n}\,e/\sqrt{\epsilon\epsilon_{0}k_{B}T}$ for monovalent ions with concentration $n$, with $e$ the fundamental charge, $\epsilon\epsilon_{0}$ the material permittivity, and $k_{B}T$ the thermal energy. This free energy $F_{\rho}$ can be expressed in Fourier space as
@@
where $\tilde{\mathbf{P}}_{\mathbf{k}}\equiv\int\mathrm{d}\mathbf{r}\,e^{-i\mathbf{k}\cdot\mathbf{r}}\mathbf{P}(\mathbf{r})$. Even in cases where $\rho(\mathbf{r})$ vanishes in the bulk, we may get contributions to $F_{\rho}$ due to charges at the boundaries. Note that our expression in Eq. (2) assumes that the screened Coulomb interaction occurs throughout the system, including outside of the domain with nonzero polarization $\mathbf{P}$.
@@
$F$ =${}_{n}=\int\mathrm{d}\mathbf{r}\left[\frac{K_{1}}{2}(\nabla\cdot\mathbf{n})^{2}+\frac{K_{2}}{2}(\mathbf{n}\cdot(\nabla\times\mathbf{n}))^{2}+\frac{K_{3}}{2}(\mathbf{n}\times(\nabla\times\mathbf{n}))^{2}\right],$ (3)
with $K_{1}$, $K_{2}$, and $K_{3}$ the splay, twist, and bend elastic constants, respectively [21]. Note that, in general, $\mathbf{P}$ and $\mathbf{n}$ have to be treated separately and are not necessarily colinear. Moreover, the proper order parameter for the nematic portion is a symmetric, traceless tensor $Q$. The more general starting point for describing the $\mathrm{N}_{\mathrm{F}}$ is given in, e.g., Ref. [22]. Here we only look at the competition between elastic distortions of the nematic order and the electrostatic self-interaction, assuming that the nematic order always aligns with $\mathbf{P}$. Any other effects will be incorporated in renormalizations of the constants (e.g., the splay $K_{1}$ and the dielectric constant $\epsilon$) and we leave a more general analysis for future work. Note also that we are neglecting the saddle-splay elastic contribution which can play a role at sample boundaries or at disclinations [23]. In our case, we will have strong anchoring conditions and any singularities will occur at polarization domain boundaries, which we will treat by introducing a phenomenological description of the energetic cost for domain wall formation, which may include effects from the saddle-splay.
We look at $\mathbf{P}$ configurations of the form:
@@
$\sum_{m=-\infty}^{\infty}A_{m}A^{*}_{m-n}=\delta_{n},$ (7)
with $\delta_{n}=1$ for $n=0$ and $\delta_{n}=0$ otherwise. Provided we have a thick sample with $h\gg\lambda_{D}$, we substitute Eqs. (4,6) into Eq. (2) and find that
@@
where $\alpha_{n}\equiv\,R\sqrt{[k_{z}(n+1)]^{2}+\kappa^{2}}$ and $I_{1}(\alpha)$, $K_{1}(\alpha)$ are the modified Bessel functions of the first and second kind, respectively. It is worth noting that $I_{1}(\alpha)K_{1}(\alpha)$ is a monotonically decreasing function of $\alpha$, meaning that the electrostatic energy favors large values of $\alpha_{n}\sim k_{z}$.
@@
which is minimized for $k_{z}\rightarrow 0$. Higher order modes $A_{n}$ with $|n|>0$ cost more elastic energy. This is less obvious for the electrostatic interaction in Eq. (8), but Khachaturyan argues on general grounds [20] that there is a stable free energy minimum with $A_{n}=0$ for all $|n|>0$.
@@
where $\alpha_{0}=R\sqrt{k_{z}^{2}+\kappa^{2}}$ and $\sum_{n}|A_{n}|^{2}=A_{0}^{2}=(A_{0}^{*})^{2}=1$ from Eq. (7). The total free energy in Eq. (11) now can be minimized with respect to $k_{z}$ for $\alpha_{0},R/\lambda_{D}\gg 1$ [so that $I_{1}(\alpha_{0})K_{1}(\alpha_{0})\approx(2\alpha_{0})^{-1}$]. We find a minimum free energy at $k_{z}=k_{z}^{*}$, which corresponds to a preferred pitch $\lambda_{z}\equiv 2\pi/k_{z}^{*}$ that reads
@@
Substituting in reasonable values $\epsilon=10$, $P_{0}=4.4\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$, $K_{2}=10~{}\mathrm{pN}$, and $R=100~{}\mu\mathrm{m}$, we find a pitch of $\lambda_{z}\approx 0.4~{}\mu\mathrm{m}$ assuming no screening $(\kappa=0$) and increasing with increasing $\kappa$. This result is consistent with the previously reported data [10] and with the observation that twists disappear in our thinnest cells [see Fig. 1a-c]. Note, however, that $\lambda_{z}$ in the considered model of an infinitely long cylinder cannot be compared directly to the experimental data obtained in Fig. 1 for flat planar samples of large area since the model does not account for the anchoring effects at the lateral surface of the cylinder. In the experiments, the azimuthal anchoring at the top and bottom plates will typically force the sample to adopt either a uniform structure or a twist by $\pi$ from the top to bottom of the cell, as discussed below.
This result for $\lambda_{z}$ was also recently found by Paik and Selinger using a different analysis [24]. Note that, due to screening, the twisted state occurs only when the concentration of ions is below some critical value: